# AFaCTA: Assisting the Annotation of Factual Claim Detection with Reliable LLM Annotators

### v1

@@ 1 Introduction
Facts entangled with Opinions |
Example 1: We are tackling other needed projects to increase capacity like six-laning I-10 in West Mobile from Theodore to Irvington. Fact part: The sentence presents a clear and explicit fact about a project. Opinion part: the project’s necessity is a subjective judgment. Example 2: We are so thankful that we haven’t suffered any loss of life, and it’s always heartening to see and hear stories of Alaskans pitching in to help each other. Fact part: no people die in the storm (according to contexts). Example 3: I thank the legislature for standing with my administration and the people of Alaska by funding this effort. Fact part: they fund the effort of resource development (according to contexts). |
Checkworthy but NOT verifiable |
Example 1: Democrats and the Media need to stop using the #Coronavirus to politicize things and scare people. It’s irresponsible. This is not the time to try and gain political points or headlines from scaring people! This tweet is labeled as check-worthy by CheckThat!-2021 (Nakov et al., 2021) since it is a polarized political opinion. However, the Democrats’ and Media’s intention is subjectively interpreted and cannot be verified by objective evidence. Example 2: Trump’s preference for well-done steaks topped with ketchup. This is an unverifiable personal preference. However, it is politicized and used to criticize political figures, thus making it checkworthy. |
Verifiable but (maybe) NOT check-worthy |
Example 1: Italy’s Prime Minister Giuseppe Conte has announced that the whole of the country is being put on lockdown in an attempt to contain the #coronavirus outbreak. This tweet with verifiable fact is labeled as NOT checkworthy by CheckThat!-2021. Example 2: Zee News: Petrol price reduced by Rs 2.69 CNN: Petrol price reduced by Rs 2.69 BBC: Petrol price reduced by Rs 2.69 NDTV: China is sending Corona Virus to the world via mails and WhatsApp. This tweet cites news with verifiable facts. But it is labeled NOT checkworthy by CheckThat!-2021. |
Context of Claims |
Example: …Those with schizophrenia spectrum and psychosis disorders, many self-medicating with drugs or alcohol addictions. That’s precisely what our encampment resolution grants and our new CARE Court seek to address. Getting people off the streets, out of tents, and into housing and treatment is essential to making our streets safe for everyone, but public safety certainly isn’t just about homelessness… This claim defines the duty of CARE but is not self-contained. It is hard to determine its verifiability without the full semantic information in context. |
@@ 2.1 Discrepancies in Prior Work
Claim conceptions: The term “claim detection” is used not only in fact-checking but also in other areas of research, for example, argument mining (Boland et al., 2022). However, this term refers to different concepts in different research areas. In fact-checking, claim detection aims at identifying objective information in statements, which can be ruled factually wrong or correct according to evidence (Thorne et al., 2018; Arslan et al., 2020; Gangi Reddy et al., 2022), and unverifiable subjective statements are usually not considered as factual claims. In contrast, in argument mining, claim detection aims at identifying the core argument or point of view referring to what is being argued about (Habernal and Gurevych, 2017). Therefore, both objective and subjective information can be identified as claims depending on their role in the discourse (Daxenberger et al., 2017; Chakrabarty et al., 2019).
@@ 4 PoliClaim Dataset
| PoliClaim | Train | |||
| PoliClaim | Train | |||
| PoliClaim | Train |
@@
The PoliClaim test set (PoliClaim) was annotated by two human experts333PhD students who are familiar with the domain of political speeches in the U.S. and COVID-related claims and have good knowledge of the literature on claim detection., who had no access to AFaCTA’s output when annotating. The experts achieved a substantial Cohen’s Kappa of 0.69 in independent annotation before the discussion. Then, they had meetings to resolve disagreements and develop gold labels. Disagreements were mainly caused by ambiguous verifiability, see Appendix A for disagreement resolving. Our annotation guideline, an instantiation of our factual claim definition, can be found in Appendix B.
@@
After verifying the performance of AFaCTA using the test sets (see more in Section 5.1), we annotated the training set with the tool’s assistance, imitating its expected use case of assisting annotation. The perfectly consistent samples were labeled directly with GPT-4 AFaCTA, while the inconsistent samples were left for human annotation. We randomly sampled 8 speeches and manually re-labeled the inconsistent annotations from AFaCTA, leading to PoliClaim where all annotations are labeled with perfect consistency or human supervision. The perfectly consistent samples in the rest of the speeches fall into PoliClaim while the inconsistent samples fall into PoliClaim. The statistics of datasets can be found in Table 2.
@@ 5 Experiments
| Agreement | Accuracy | Agreement | Accuracy | Agreement | Accuracy | |
| GPT-3.5 | ||||||
| GPT-4 | ||||||
| Experts | / | / | / | / | ||
@@ 5.1 AFaCTA Annotation Performance
Takeaway: With AFaCTA’s self-consistency calibration, auto-annotation of perfectly consistent samples can be reliably adopted to reduce manual effort (also see Section 5.5). In the case of PoliClaim, only 51.22% needs further supervision, while 48.78% of manual effort is saved with GPT-4-AFaCTA.
@@ 5.4 Domain Agnostic AFaCTA
The reasoning logic of AFaCTA is not restricted to the political speech domain. To verify its performance on the social media domain, we conduct the analyses in Section 5.1 and Section 5.3 again on the CheckThat!-2021 (Nakov et al., 2021) development set. Experiment results are similar to those on PoliClaim (see Appendix J). Therefore, AFaCTA may assist factual claim annotation in various domains.
@@ 5.5 AFaCTA Delivers Useful Annotations
To explore whether AFaCTA’s annotation can replace or augment manual annotation in training classifiers, we train hundreds of classifiers with different combinations of PoliClaim (AFaCTA annotations + Human Supervision), PoliClaim (AFaCTA perfectly consistent annotations), and PoliClaim (AFaCTA inconsistent annotations). All results are averaged over random seeds of 42, 43, and 44, and are supported with statistical significance tests (see Appendix L). 444This section presents RoBERTa (Liu et al., 2019) results. Appendix M presents similar DistilBERT (Sanh et al., 2019) results as side findings. Detailed fine-tuning settings are in Appendix K.
Using only gold, silver, or bronze data: We first gradually increase the number of training data points (by 100 per step) of the same quality. Results are shown in Figure 3. We observe the same phenomenon as previous work (Stammbach et al., 2023b) where the marginal accuracy gain drops while adding more data. The PoliClaim and PoliClaim curves roughly follow the same growing trend, approaching GPT-4’s aggregated performance. This indicates that the perfectly consistent annotations (silver) from AFaCTA can strongly substitute for manually annotated data. The PoliClaim curve is slightly higher, showing that learning from human-supervised hard samples (inconsistent annotations of AFaCTA) is beneficial. The PoliClaim curve is much lower, showing that the noisy, inconsistent annotations harm the classifier training.
Augmenting training with auto-annotated data: When the manual annotation budget is limited, can we augment the dataset with automatic annotation? In Figure 4, we gradually augment the PoliClaim data with automatically annotated ones (100 per step). It can be observed that: (1) The performance increases more with PoliClaim data augmentation, showing that the data quality is important in data augmentation. (2) Compared to augmenting the full PoliClaim dataset, augmentation results in more improvement when there are only 500 PoliClaim data. Therefore, high-quality automatic annotation is more helpful when the manual annotation budget is limited. (3) Combining gold and silver data leads to classifiers that outperform aggregated GPT-4 reasoning, demonstrating that extending training data with LLM annotation is a promising approach to achieving better performance.
@@ 7 Conclusion
Check-Worthiness. The objective of factual claim detection is to prioritize claims that are both verifiable and check-worthy, maximizing the use of potentially limited fact-checking resources. However, in this project, we focus on verifiability without exploiting the other important aspect: checkworthiness. Konstantinovskiy et al. (2020) argues that the definition of check-worthiness is subjective. However, it is possible to define a claim’s checkworthiness according to its context. For example, is the claimer an influential person or media? Is the topic controversial? There has already been work that takes some contextual information (e.g., claimer, topic, etc.) into account (Gangi Reddy et al., 2022). In future work, we will explore deterministic and efficient ways to define and annotate checkworthiness leveraging rich contextual information.
@@
Limited Expert Annotators. We only evaluate AFaCTA’s annotation performance against two experts, which may lead to potential bias. We fail to hire more expert annotators mainly because expert annotation is extremely expensive, and it is hard to find more experts with good knowledge about factual claim definitions. As compensation, we release all expert annotations and detailed error analyses where the potential bias can be analyzed. Besides, adding unsupervised LLM-annotated data continuously improves the accuracy on PoliClaim, demonstrating that our human labeling on PoliClaim has very limited bias.
Only GPT-4 Is Reliable. We find that only GPT-4-AFaCTA outperforms human experts on perfectly consistent samples. GPT-3.5 achieves promising results but tends to produce false negative errors. Although GPT-4 is much cheaper than human supervision, it is close-sourced and is comparatively more expensive than other LLMs. In future work, we will study how to use open-sourced models to produce high-quality annotations. Specifically, we will explore (1) training the model to better understand the annotation guideline; (2) leveraging internal certainties like output logits; and (3) extending the spectrum of self-consistency levels with cheaper inference.
@@ Acknowledgements
This paper has received funding from the Swiss National Science Foundation (SNSF) under the project ‘How sustainable is sustainable finance? Impact evaluation and automated greenwashing detection’ (Grant Agreement No. 100018_207800).
@@ Appendix F AFaCTA with Open-sourced LLMs
| PoliClaim | CheckThat!2021-dev | |||||
| Agreement | Accuracy | Consistency | Agreement | Accuracy | Consistency | |
| zephyr-7b- | 0.205 | 66.18 | 0.49 | 0.539 | 77.86 | 5.00 |
| llama-2-13b-chat | 0.306 | 56.74 | 0.00 | 0.260 | 50.71 | 1.43 |
| GPT-3.5 | 0.510 | 76.74 | 43.38 | 0.359 | 69.29 | 44.29 |
| GPT-4 | 0.615 | 86.27 | 48.78 | 0.437 | 86.43 | 57.85 |
@@ Appendix H Performance of Each AFaCTA Step
| Step 1 | Step 2 | Step 3 | ||||
| Agreement | Accuracy | Agreement | Accuracy | Agreement | Accuracy | |
| GPT-3.5 | 0.458 | 73.16 | 0.452 | 78.06 | 0.546 | 66.42 |
| GPT-4 | 0.633 | 85.54 | 0.437 | 79.90 | 0.630 | 73.28 |
@@ Appendix I Self-Consistency CoT
| Agreement | Accuracy | Agreement | Accuracy | Agreement | Accuracy | |
| GPT-3.5 | ||||||
| GPT-4 | ||||||
| Experts | / | / | / | / | ||
@@
Similar observations as PoliClaim can be drawn. GPT-4 AFaCTA outperforms experts on perfectly consistent samples and underperforms on inconsistent samples. GPT-3.5 also achieves a moderate agreement with human experts on perfectly consistent samples. Error analysis shows that GPT-3.5’s error concentrates on false negatives, similar to its behavior in the political speech domain (see Table 12).
@@
We conduct a statistical significance test to show that different training set combinations of PoliClaim, PoliClaim, and PoliClaim lead to statistically significant differences in fine-tuning claim detectors. We first conduct a Student-t test for each training combination based on the results of three random seeds and then aggregate p-values using Fisher’s method. For example, to compare “only PoliClaim” vs. only “PoliClaim”, we use the following formula:
@@
| Only S | < | Only G | 5.54e-3 | 8.89e-5 |
| Only B | < | Only S | 2.39e-36 | 5.79e-51 |
| Only B | < | Only G | 1.88e-20 | 6.03e-29 |
| 500 G + B | < | 500 G + S | 1.50e-28 | 1.82e-30 |
| 1000 G + B | < | 1000 G + S | 8.13e-13 | 3.30e-8 |
| 1500 G + B | < | 1500 G + S | 2.19e-16 | 1.69e-15 |
| All G + B | < | All G + S | 3.68e-9 | 1.36e-13 |
@@
We conduct the data augmentation experiments in Section 5.5 on both RoBERTa (Figure 8) and DistilBERT (Figure 9) with a different number of PoliClaim data (500, 1000, 1500, and 1936). Similar conclusions as Section 5.5 can be drawn: perfectly consistent (silver) data are better at augmentation than inconsistent (bronze) data. Figure 10 also shows a clear trend. When the manual annotation budget is more restricted, more augmentation data are needed to achieve a comparable performance.
@@
False Positive: over-sensitive to granular, unspecific, or not-explicitly-presented facts |
Error 1: I just want to thank you, thank you members of the Legislature for all you did these past two years to keep us safe. Error reason: recognizing ’members of the legislature did some thing’ as fact, which is too vague. Error 2: It’s true from the Flatirons to Fishers Peak to Pikes Peak to Longs Peak and beyond. Error reason: location names are recognized as facts. Error 3: Sheriff Pelle, firefighters, and emergency responders, please stand so we can thank you for the lifesaving work that you do every day. Error reason: people’s appearance at the event is recognized as a fact, which is not explicitly presented. Error 4: I’m glad to be back at the capitol addressing the Legislature in person, and I thank you for the invitation to speak to you tonight. Error reason: identify the speaker’s back and addressing legislature as facts. |
False Negative: not enough context |
Error 5: It’s the result of great investment decisions, policies, vision, and direction. Error reason: “it” here refers to the return of pension fund, which is in a far context. But AFaCTA only considers a one-sentence context. Error 6: Alaskans won’t accept that we can’t get anything done because it’s an election year. Error reason: it claims a fact that this year is an election year, but the model comprehends this as a hypothetical condition, due to its lack of context that 2022 is the election year for Alaska. |
@@
False Positive: over-sensitive to unspecific facts |
@@
Error 1: The fresh mountain air that so many people associate with Colorado isn’t a given. Error reason: identifying “the fresh air is not a given” as a fact, which is unspecific and leans towards unverifiable opinion. |
False Negative: not enough specific detail or context and thus not verifiable |

### v3

@@ 1 Introduction
| Facts entangled with Opinions | Example 1: We are tackling other needed projects to increase capacity like six-laning I-10 in West Mobile from Theodore to Irvington. Fact part: The sentence presents a clear and explicit fact about a project. Opinion part: the project’s necessity is a subjective judgment. Example 2: We are so thankful that we haven’t suffered any loss of life, and it’s always heartening to see and hear stories of Alaskans pitching in to help each other. Fact part: no people die in the storm (according to contexts). Example 3: I thank the legislature for standing with my administration and the people of Alaska by funding this effort. Fact part: they fund the effort of resource development (according to contexts). |
| Checkworthy but NOT verifiable | Example 1: Democrats and the Media need to stop using the #Coronavirus to politicize things and scare people. It’s irresponsible. This is not the time to try and gain political points or headlines from scaring people! This tweet is labeled as check-worthy by CheckThat!-2021 (Nakov et al., 2021) since it is a polarized political opinion. However, the Democrats’ and Media’s intention is subjectively interpreted and cannot be verified by objective evidence. Example 2: Trump’s preference for well-done steaks topped with ketchup. This is an unverifiable personal preference. However, it is politicized and used to criticize political figures, thus making it checkworthy. |
| Verifiable but (maybe) NOT check-worthy | Example 1: Italy’s Prime Minister Giuseppe Conte has announced that the whole of the country is being put on lockdown in an attempt to contain the #coronavirus outbreak. This tweet with verifiable fact is labeled as NOT checkworthy by CheckThat!-2021. Example 2: Zee News: Petrol price reduced by Rs 2.69 CNN: Petrol price reduced by Rs 2.69 BBC: Petrol price reduced by Rs 2.69 NDTV: China is sending Corona Virus to the world via mails and WhatsApp. This tweet cites news with verifiable facts. But it is labeled NOT checkworthy by CheckThat!-2021. |
| Context of Claims | Example: …Those with schizophrenia spectrum and psychosis disorders, many self-medicating with drugs or alcohol addictions. That’s precisely what our encampment resolution grants and our new CARE Court seek to address. Getting people off the streets, out of tents, and into housing and treatment is essential to making our streets safe for everyone, but public safety certainly isn’t just about homelessness… This claim defines the duty of CARE but is not self-contained. It is hard to determine its verifiability without the full semantic information in context. |
@@ 2.1 Discrepancies in Prior Work
Claim conceptions: The term “claim detection” is used not only in fact-checking but also in other areas of research, for example, argument mining (Boland et al., 2022). However, this term refers to different concepts in different research areas. In fact-checking, claim detection aims at identifying objective information in statements, which can be ruled factually wrong or correct according to evidence (Thorne et al., 2018; Arslan et al., 2020; Gangi Reddy et al., 2022), and unverifiable subjective statements are usually not considered as factual claims. In contrast, in argument mining, claim detection aims at identifying the core argument or point of view referring to what is being argued about (Habernal and Gurevych, 2017). Therefore, both objective and subjective information can be identified as claims depending on their role in the discourse (Daxenberger et al., 2017; Chakrabarty et al., 2019). The intermixing of such concepts has led to dataset misuse issues in research: for instance, Gupta et al. (2021) annotate a claim detection dataset for fack-checking COVID-19 tweets. However, the dataset is jointly trained and evaluated with claim detection datasets for argument mining (Peldszus and Stede, 2015; Stab and Gurevych, 2017, inter alia), which potentially harms the soundness of the results.
@@ 4 PoliClaim Dataset
| PoliClaimgold | Train | |||
| PoliClaimsilver | Train | |||
| PoliClaimbronze | Train |
@@
The PoliClaim test set (PoliClaimtest) was annotated by two human experts333PhD students who are familiar with the domain of political speeches in the U.S. and COVID-related claims and have good knowledge of the literature on claim detection., who had no access to AFaCTA’s output when annotating. The experts achieved a substantial Cohen’s Kappa of 0.69 in independent annotation before the discussion. Then, they had meetings to resolve disagreements and develop gold labels. Disagreements were mainly caused by ambiguous verifiability, see Appendix A for disagreement resolving. Our annotation guideline, an instantiation of our factual claim definition, can be found in Appendix B.
@@
After verifying the performance of AFaCTA using the test sets (see more in Section 5.1), we annotated the training set with the tool’s assistance, imitating its expected use case of assisting annotation. The perfectly consistent samples were labeled directly with GPT-4 AFaCTA, while the inconsistent samples were left for human annotation. We randomly sampled 8 speeches and manually re-labeled the inconsistent annotations from AFaCTA, leading to PoliClaimgold where all annotations are labeled with perfect consistency or human supervision. The perfectly consistent samples in the rest of the speeches fall into PoliClaimsilver while the inconsistent samples fall into PoliClaimbronze. The statistics of datasets can be found in Table 2.
@@ 5 Experiments
| Agreement | Accuracy | Agreement | Accuracy | Agreement | Accuracy | |||
| GPT-3.5 | ||||||||
| GPT-4 | ||||||||
| Experts | / | / | / | / | ||||
@@ 5.1 AFaCTA Annotation Performance
Takeaway: With AFaCTA’s self-consistency calibration, auto-annotation of perfectly consistent samples can be reliably adopted to reduce manual effort (also see Section 5.5). In the case of PoliClaimtest, only 51.22% needs further supervision, while 48.78% of manual effort is saved with GPT-4-AFaCTA.
@@ 5.4 Domain Agnostic AFaCTA
The reasoning logic of AFaCTA is not restricted to the political speech domain. To verify its performance on the social media domain, we conduct the analyses in Section 5.1 and Section 5.3 again on the CheckThat!-2021 (Nakov et al., 2021) development set. Experiment results are similar to those on PoliClaimtest (see Appendix J). Therefore, AFaCTA may assist factual claim annotation in various domains.
@@ 5.5 AFaCTA Delivers Useful Annotations
To explore whether AFaCTA’s annotation can replace or augment manual annotation in training classifiers, we train hundreds of classifiers with different combinations of PoliClaimgold (AFaCTA annotations + Human Supervision), PoliClaimsilver (AFaCTA perfectly consistent annotations), and PoliClaimbronze (AFaCTA inconsistent annotations). All results are averaged over random seeds of 42, 43, and 44, and are supported with statistical significance tests (see Appendix L). 444This section presents RoBERTa (Liu et al., 2019) results. Appendix M presents similar DistilBERT (Sanh et al., 2019) results as side findings. Detailed fine-tuning settings are in Appendix K.
Using only gold, silver, or bronze data: We first gradually increase the number of training data points (by 100 per step) of the same quality. Results are shown in Figure 3. We observe the same phenomenon as previous work (Stammbach et al., 2023b) where the marginal accuracy gain drops while adding more data. The PoliClaimgold and PoliClaimsilver curves roughly follow the same growing trend, approaching GPT-4’s aggregated performance. This indicates that the perfectly consistent annotations (silver) from AFaCTA can strongly substitute for manually annotated data. The PoliClaimgold curve is slightly higher, showing that learning from human-supervised hard samples (inconsistent annotations of AFaCTA) is beneficial. The PoliClaimbronze curve is much lower, showing that the noisy, inconsistent annotations harm the classifier training.
Augmenting training with auto-annotated data: When the manual annotation budget is limited, can we augment the dataset with automatic annotation? In Figure 4, we gradually augment the PoliClaimgold data with automatically annotated ones (100 per step). It can be observed that: (1) The performance increases more with PoliClaimsilver data augmentation, showing that the data quality is important in data augmentation. (2) Compared to augmenting the full PoliClaimgold dataset, augmentation results in more improvement when there are only 500 PoliClaimgold data. Therefore, high-quality automatic annotation is more helpful when the manual annotation budget is limited. (3) Combining gold and silver data leads to classifiers that outperform aggregated GPT-4 reasoning, demonstrating that extending training data with LLM annotation is a promising approach to achieving better performance. One of the best RoBERTa checkpoints trained on all PoliClaimgold and PoliClaimsilver is available on HuggingFace555https://huggingface.co/JingweiNi/roberta-base-afacta.
@@ 6 Related Work
7 Discussions
7.1 Check-Worthiness
The objective of factual claim detection is to prioritize claims that are both verifiable and check-worthy, maximizing the use of potentially limited fact-checking resources. However, in this project, we focus on verifiability without exploiting the other important aspect: checkworthiness. Konstantinovskiy et al. (2020) argues that the definition of check-worthiness is subjective. However, it is possible to define a claim’s checkworthiness according to its context. For example, is the claimer an influential person or media? Is the topic controversial? There has already been work that takes some contextual information (e.g., claimer, topic, etc.) into account (Gangi Reddy et al., 2022). Future work may explore deterministic and efficient ways to define and annotate checkworthiness leveraging rich contextual information.
7.2 Only GPT-4 Is Reliable
We find that only GPT-4-AFaCTA outperforms human experts on perfectly consistent samples. GPT-3.5 achieves promising results but tends to produce false negative errors. Although GPT-4 is much cheaper than human supervision, it is close-sourced and is comparatively more expensive than other LLMs. Future work may study how to use open-sourced models to produce high-quality annotations. Specifically, future work may explore (1) training the model to better understand the annotation guideline; (2) leveraging internal certainties like output logits; and (3) extending the spectrum of self-consistency levels with cheaper inference.
@@ 8 Conclusion
Besides, AFaCTA step 2 and 3 cost (approximately) 6.5x and 8.5x more tokens than step 1. Although step 2 and 3 bring self-consistency calibration and performance gain through aggregation, the marginal benefit of API cost is far from perfect.
@@
Limited Expert Annotators. We only evaluate AFaCTA’s annotation performance against two experts, which may lead to potential bias. We fail to hire more expert annotators mainly because expert annotation is extremely expensive, and it is hard to find more experts with good knowledge about factual claim definitions. As compensation, we release all expert annotations and detailed error analyses where the potential bias can be analyzed. Besides, adding unsupervised LLM-annotated data continuously improves the accuracy on PoliClaimtest, demonstrating that our human labeling on PoliClaimtest has very limited bias.
@@ Acknowledgements
This paper has received funding from the Swiss National Science Foundation (SNSF) under the project ‘How sustainable is sustainable finance? Impact evaluation and automated greenwashing detection’ (Grant Agreement No. 100018_207800). It is also funded by grant from Hasler Stiftung for the Research Program Responsible AI with the project “Scientific Claim Verification.”
@@ Appendix F AFaCTA with Open-sourced LLMs
| PoliClaimtest | CheckThat!2021-dev | ||||||
| Agreement | Accuracy | Consistency | Agreement | Accuracy | Consistency | ||
| zephyr-7b- | 0.205 | 66.18 | 0.49 | 0.539 | 77.86 | 5.00 | |
| llama-2-13b-chat | 0.306 | 56.74 | 0.00 | 0.260 | 50.71 | 1.43 | |
| GPT-3.5 | 0.510 | 76.74 | 43.38 | 0.359 | 69.29 | 44.29 | |
| GPT-4 | 0.615 | 86.27 | 48.78 | 0.437 | 86.43 | 57.85 | |
@@ Appendix H Performance of Each AFaCTA Step
| Step 1 | Step 2 | Step 3 | ||||||
| Agreement | Accuracy | Agreement | Accuracy | Agreement | Accuracy | |||
| GPT-3.5 | 0.458 | 73.16 | 0.452 | 78.06 | 0.546 | 66.42 | ||
| GPT-4 | 0.633 | 85.54 | 0.437 | 79.90 | 0.630 | 73.28 | ||
@@ Appendix I Self-Consistency CoT
| Agreement | Accuracy | Agreement | Accuracy | Agreement | Accuracy | |||
| GPT-3.5 | ||||||||
| GPT-4 | ||||||||
| Experts | / | / | / | / | ||||
@@
Similar observations as PoliClaimtest can be drawn. GPT-4 AFaCTA outperforms experts on perfectly consistent samples and underperforms on inconsistent samples. GPT-3.5 also achieves a moderate agreement with human experts on perfectly consistent samples. Error analysis shows that GPT-3.5’s error concentrates on false negatives, similar to its behavior in the political speech domain (see Table 12).
@@
We conduct a statistical significance test to show that different training set combinations of PoliClaimgold, PoliClaimsilver, and PoliClaimbronze lead to statistically significant differences in fine-tuning claim detectors. We first conduct a Student-t test for each training combination based on the results of three random seeds and then aggregate p-values using Fisher’s method. For example, to compare “only PoliClaimgold” vs. only “PoliClaimsilver”, we use the following formula:
@@
| Only S | < | Only G | 5.54e-3∗ | 8.89e-5∗∗ |
| Only B | < | Only S | 2.39e-36∗∗ | 5.79e-51∗∗ |
| Only B | < | Only G | 1.88e-20∗∗ | 6.03e-29∗∗ |
| 500 G + B | < | 500 G + S | 1.50e-28∗∗ | 1.82e-30∗∗ |
| 1000 G + B | < | 1000 G + S | 8.13e-13∗∗ | 3.30e-8∗∗ |
| 1500 G + B | < | 1500 G + S | 2.19e-16∗∗ | 1.69e-15∗∗ |
| All G + B | < | All G + S | 3.68e-9∗∗ | 1.36e-13∗∗ |
@@
We conduct the data augmentation experiments in Section 5.5 on both RoBERTa (Figure 8) and DistilBERT (Figure 9) with a different number of PoliClaimgold data (500, 1000, 1500, and 1936). Similar conclusions as Section 5.5 can be drawn: perfectly consistent (silver) data are better at augmentation than inconsistent (bronze) data. Figure 10 also shows a clear trend. When the manual annotation budget is more restricted, more augmentation data are needed to achieve a comparable performance.
@@
| False Positive: over-sensitive to granular, unspecific, or not-explicitly-presented facts | Error 1: I just want to thank you, thank you members of the Legislature for all you did these past two years to keep us safe. Error reason: recognizing ’members of the legislature did some thing’ as fact, which is too vague. Error 2: It’s true from the Flatirons to Fishers Peak to Pikes Peak to Longs Peak and beyond. Error reason: location names are recognized as facts. Error 3: Sheriff Pelle, firefighters, and emergency responders, please stand so we can thank you for the lifesaving work that you do every day. Error reason: people’s appearance at the event is recognized as a fact, which is not explicitly presented. Error 4: I’m glad to be back at the capitol addressing the Legislature in person, and I thank you for the invitation to speak to you tonight. Error reason: identify the speaker’s back and addressing legislature as facts. |
| False Negative: not enough context | Error 5: It’s the result of great investment decisions, policies, vision, and direction. Error reason: “it” here refers to the return of pension fund, which is in a far context. But AFaCTA only considers a one-sentence context. Error 6: Alaskans won’t accept that we can’t get anything done because it’s an election year. Error reason: it claims a fact that this year is an election year, but the model comprehends this as a hypothetical condition, due to its lack of context that 2022 is the election year for Alaska. |
@@
| False Positive: over-sensitive to unspecific facts | Error 1: The fresh mountain air that so many people associate with Colorado isn’t a given. Error reason: identifying “the fresh air is not a given” as a fact, which is unspecific and leans towards unverifiable opinion. |